# Simulating many-engine spacecraft: Exceeding 1 quadrillion degrees of freedom via information geometric regularization

### v3

@@ 48 $(\rho,\rho\mathbf{u},E,\alpha)\leftarrow(\rho,\rho\mathbf{u},E,\alpha)+\mathrm{d}t\cdot\mathtt{rhs}$
VI How Performance was Measured
VI-A HPC platforms
@@
We performed calculations on LLNL El Capitan, OLCF Frontier, and CSCS Alps. Table˜II shows the performance attributes of these systems.
VI-A1 LLNL El Capitan
@@
The El Capitan nodes have four AMD MI300A APUs. Each APU combines 24 EPYC Zen 4 CPU cores and 228 CDNA 3 compute units with a single $128\text{\,}\mathrm{GB}\text{/}$ layer of HBM3 memory for both processor types. Nodes are interconnected via HPE Cray Slingshot-11 Ethernet in a dragonfly topology with four $200\text{\,}\mathrm{GB}\text{/}\mathrm{s}$ NICs per node. The GPUs can be programmed with unified shared memory (USM) to avoid duplicate host–device memory addresses. We use this strategy via OpenMP target offload (see section˜V-E).
VI-A2 OLCF Frontier
@@
Each Frontier node has one 64-core AMD EPYC Trento CPU and four AMD MI250X GPUs. Each MI250X GPU contains two Graphics Compute Dies (GCDs) with $64\text{\,}\mathrm{GB}\text{/}$ of HBM2E each (or $128\text{\,}\mathrm{GB}\text{/}$ per GPU). The total system memory is $9.6\text{\,}\mathrm{PB}\text{/}$, equally split between HBM2E GPU and DDR4 CPU memory. Frontier uses HPE Slingshot Ethernet for interconnects with a 3-hop Dragonfly topology and four $200\text{\,}\mathrm{GB}\text{/}\mathrm{s}$ NICs per node attached to the GPUs.
VI-A3 CSCS Alps
@@
Alps nodes have 4 NVIDIA Grace Hopper GH200 superchips. Each GH200 has a Hopper GPU with $4\text{\,}\mathrm{TB}\text{/}\mathrm{s}$ of memory bandwidth for its $96\text{\,}\mathrm{GB}\text{/}$ of HBM3 memory and a Grace CPU with 72 Arm-v9 cores with $500\text{\,}\mathrm{GB}\text{/}\mathrm{s}$ bandwidth for its $120\text{\,}\mathrm{GB}\text{/}$ of LPDDR5 memory. NVLink-C2C connects the CPU and GPU, enabling efficient data movement. We use this new capability herein (see section˜V-E). Alps’s nodes interconnect through HPE Slingshot with $200\text{\,}\mathrm{GB}\text{/}\mathrm{s}$ injection bandwidth per superchip.
VI-B Software environment and performance baseline
@@
We base our implementation on MFC [23, 22], a compressible flow solver that implements state-of-the-art numerical shock capturing and the IGR implementation of this work. All of MFC’s schemes scale ideally to 100% of LLNL El Capitan, OLCF Frontier, and CSCS Alps, among other previous and existing flagship supercomputers [24, 25].
@@
MFC offloads computation onto GPU and superchip/APU devices via either OpenMP or OpenACC, and uses metaprogramming to abstract away and automate vendor-specific or otherwise burdensome optimizations. MFC has a history of being used to simulate compressible multi-species, phase, and chemically reacting fluid flows [26, 27, 28, 29]. Performance results are measured using a representative three-dimensional simulation of the exhaust plume of a single Mach 10 jet.
@@
We use MFC’s optimized implementation of WENO nonlinear reconstructions and HLLC approximate Riemann solves as a baseline for performance comparisons [22]. This implementation matches or outperforms other codebases with the same and similar reconstruction schemes. On Frontier and El Capitan we test HPE’s CCE 19.0.0 compiler and a pre-release of AMD’s Next Generation Flang compiler with HPE Cray MPICH-8.1.31 for messaging. For both compilers, we specify -O3 for optimization. With CCE, we also apply the -haggress flag to improve performance in large kernels. On Alps, we use a pre-release of the NVIDIA HPC SDK 25.9 and HPE Cray MPICH 8.1.30 for messaging. We also use the compile-time flag -gpu=fastmath, yielding performance improvement for single precision computation on NVIDIA devices. The latest NVIDIA HPC SDK exhibits an FP64 performance regression, instead using NVHPC SDK 24.3.
VI-C Measurement tools
@@
We measure execution time with application internal timers using standard cpu_time and system_clock procedures. On Frontier and El Capitan, we periodically sample the instantaneous GPU or APU power draw by reading in the virtual filesystem for the AMD driver. The power reading is the same as reported by rocm-smi, but reading from the virtual filesystem has lower overhead. On Frontier, rocm-smi collects GPU and HBM power only. On El Capitan rocm-smi includes CPU, GPU, and memory power. On Alps, nvidia-smi records the module (CPU, GPU, and memory) power draw.
@@
On all platforms, we post-process the results such that we only account for power draw during time-stepping, which is then averaged and multiplied by the average time per time step. The number of grid points further normalizes energy use.
VII Performance Results
VII-A Time step cost
Table˜III shows the normalized grind times for a problem solved using WENO reconstructions and HLLC approximate Riemann solves with in-core computation (current state of the art, optimized implementation in MFC [22]), IGR with in-core computation, and IGR with unified memory on one GH200 on Alps, one MI250X GCD on Frontier, and one MI300A APU on El Capitan. The grind time is defined as nanoseconds per grid cell per time step to normalize against the different problem sizes that fit within device memory. Smaller grind times indicate shorter time to solution.
GH200 metrics are collected using separate memory mode for in-core cases and unified memory mode for the out-of-core/unified cases. Performance impacts of less than 5% are observed when moving from in-core to unified computation on the GH200 architecture. For GH200 FP64 cases, we used the NVHPC 24.3 compiler, as there was a regression when switching from separate memory to unified memory in-core computation. We observed a $9\%$ performance hit for that case (see table˜III). The performance difference between in- and out-of-core unified memory builds is less than $5\%$.
@@ VII-A Time step cost
TABLE III: Wall time for representative simulations, quantified via nanoseconds per grid cell per time step. The baseline method is compared to the current work. For the AMD MI300A, USM mode is used, so no in-core quantities are presented; other unified cases use UVM. AMD device cases were run using Cray CCE 19.0.0 and AMD Flang Preview 7.0.5 compilers, using OpenACC and OpenMP offloading, with the best result shown. AMD Flang is the only current compiler option for native FP16/32 on all AMD GPU/APU devices. NVHPC is used for the GH200 cases.
@@
The increased grind time in calculations that use unified memory to increase problem size per device on Frontier and Alps results from the exchange of conservative variable buffers between the CPU and GPU at each Runge–Kutta update. On all devices, time to solution is reduced by a factor of about 4 when going from WENO to IGR in FP64. FP64 is broadly recommended for ENO-type shock-capturing due to catastrophic cancellation [15]. Our approach can even handle even mixed FP16/FP32 precision. This reduces the time to solution by a factor of at least 6 compared to the baseline.
@@ VII-B Scaling
We exceed 200T grid cells, or 1 quadrillion degrees of freedom (5 state variables) on Frontier using 37.6K MI250X GCDs (9408 nodes), accommodating $1386^{3}$ grid points per GCD with UVM, and FP16/32 mixed precision. On Alps, a problem size of $1611^{3}$ is used per GH200 and UVM with FP16/32 mixed precision, amounting to 45T grid point simulation on the full system (2688 nodes). On the recently deployed JSC JUPITER, this amounts to 100.3T grid points or 501T degrees of freedom, given its size and matching architecture to Alps. On El Capitan, we use $1380^{3}$ grid points per MI300A GPU and reach 113T grid points using 10750 nodes (11.1K nodes full system). The lower grid count on Alps and El Capitan owes purely to system size.
Figure 6: Weak scaling performance for a representative thruster problem on LLNL El Capitan (number of MI300As), OLCF Frontier (number of MI250Xs), and CSCS Alps (number of GH200s). The full system is indicated for each case. All runs use unified memory, FP16/32 mixed precision, and a 16 node configuration for scaling comparison, which ensures that all MPI communication directions are touched. On El Capitan, we see $97\%$ efficiency out to 10750 MI300As. On Alps, ideal ($\approx 100\%$) scaling is observed at 9.2K GH200s. On Frontier, we also achieve ideal ($\approx 100\%$) efficiency at 37.6K MI250X GPUs, with a maximum problem size of 200T grid points.
Strong scaling results on Alps, Frontier, and El Capitan are shown in fig.˜7. The base case uses $8$ nodes ($32$ GPUs), leading to $32$ ranks on El Capitan and Alps and $64$ ranks on Frontier, arranged in a rectilinear configuration. We use FP16/32 mixed precision for storage/computation; unified memory is enabled via UVM on Frontier and Alps and USM on El Capitan; GPU-aware MPI is used on Alps. For a 32-fold increase in device count, we achieve strong scaling efficiencies of 90%, 90%, and 86% on El Capitan, Frontier, and Alps. When increasing to the full systems, we achieve 44% (El Capitan), 44% (Frontier), and 80% (Alps) strong scaling efficiencies. Thus, one can execute an 8 node computation on the full system, decreasing time to solution by a factor of about 500.
@@
Figure 7: Strong scaling on all systems for the same problem configuration as fig.˜6. All results are for FP16/32 mixed precision, though FP32 and FP64 show similar results (FP32 shown in fig.˜8). The full system is shown for each case. Following fig.˜6, we base all speedups on an 8 node configuration, such that all communication directions are touched.
Figure 8: Strong scaling following fig.˜7, here for FP32 on Frontier. We compare the performance of the current work to the optimized baseline method as labeled. The current work uses and can accommodate 10.5B grid points per node, and the baseline accommodates 421M grid points, both of which are used for the 8-node speedup reference.
VII-C Energy efficiency
Table˜IV shows the measured energy consumption per grid cell per time step, comparing our IGR implementation to a prior-state-of-the-art WENO implementation. The problem size is adjusted to exhaust the total GPU/APU memory on a single node in double precision. On the NIVIDA GH200 we measured the in-core implementation. On the AMD MI250X and MI300A, we use the CCE 19.0.0 compiler with GPU-only (in-core) and unified memory builds.

### v4

@@ 48 $(\rho,\rho\mathbf{u},E,\alpha)\leftarrow(\rho,\rho\mathbf{u},E,\alpha)+\mathrm{d}t\cdot\mathtt{rhs}$
5.4. Optimizations
Our implementation eliminates the storage of the reconstructed states (algorithms 1 to 1), velocity gradients (algorithms 1 to 1), and fluxes (algorithms˜1, 1 and 1) and keeps all operations in a single kernel (algorithm˜1). The memory footprint is reduced by storing the intermediate variables as thread-local temporary arrays within this kernel. The algorithm only requires storing $2$ copies of the conservative variables, the net flux at each grid point, and the solution and left side of the elliptic PDE in eq.˜9.
Each thread solves an approximate Riemann problem at the grid cell–cell interface and accumulates its contribution to the right-hand side at overlapping cells via atomic operations to prevent race conditions. During the reconstruction along the 1st coordinate dimension ($x$, here), the contributions to the left-hand side of equation eq.˜9 are computed using the velocity gradients for the viscous fluxes (algorithm˜1). The elliptic PDE in eq.˜9 is solved after its left side is computed (algorithm˜1). The entropic pressure added directly to the flux in the 2nd and 3rd dimensions (algorithms 1 to 1). The flux contribution of the entropic pressure in the first dimension is completed separately after the elliptic solve (algorithms 1 to 1). The above decreases memory use $25$-fold compared to an optimized 5th-order accurate WENO and Riemann solver implementation in the same codebase (wilfong252).
Table 2. Node and full system properties of the supercomputers tested in this work. TOP500 rankings from June 2025.
Refer to caption :
Figure 4. Schematic of the chip-to-chip (C2C) transfers of intermediate time-step variables between the on-node CPU and GPU devices. The time sub-steps are $q_{1,2}$ and the full step integration is stored in $q_{1}$.
5.5. Unified Memory
The unified memory approach uses unified-shared-memory (USM) mode on the MI300A (El Capitan), AMD InfinityFabric for CPU–GPU connections on Frontier, and the NVLink C2C connection on Grace Hopper (Alps), Figure˜4 shows this strategy, using the full capacity of the compute node while expanding beyond the GPU memory without incurring a meaningful performance penalty. This allows us to store the intermediate Runge-Kutta stage on the host, reducing GPU memory use by up to a factor of $12/17$.
5.5.1. AMD approach for El Capitan
On the AMD MI300A, the CPU and GPU share a single physical HBM pool. We compile MFC with OpenMP offloading directives using AMD’s next-generation Flang compiler and OpenMP’s USM mode. All variables have a single copy in memory and are accessible from CPUs and accelerators with no data movement. CCE’s OpenACC implementation is also used for speed comparisons.
5.5.2. AMD approach for Frontier
Our approach to Frontier involves using OpenACC or OpenMP, as well as HPE’s CCE or the AMD Flang next-generation compiler, for performance comparisons between compilers and unified memory implementations.
With OpenMP, the mapping of the Fortran pointers associated with the device-resident memory within the MFC data structures is omitted. No redundant memory copy is created due to MP’s treatment of the derived type mapping with Fortran pointer components. MFC can also run on Frontier in USM mode, but was found to be less performant on Frontier than a UVM strategy discussed in section˜5.5.3 below, and is not used in this work. In summary, we reduce memory usage on both the host and device without compromising performance.
5.5.3. NVIDIA approach for Alps
The time step updates in the Runge–Kutta scheme are rearranged so that only the current sub-step is passed to the right-hand side routine. The buffer holding the previous state is used to update the current Runge–Kutta state. With this rearrangement, we only store one sub-step and the right-hand side buffer in GPU memory. Thus, the intermediary sub-step is always in CPU memory, freeing GPU memory and increasing simulation size without sacrificing performance. The time step updates access Runge–Kutta sub-steps from their physical locations via zero-copy, allowing simultaneous access to data from both GPU and CPU memory through the C2C NVLink. We include flexibility in hosting the IGR temporary variables in either GPU or CPU memory, further reducing the memory footprint from $12/17$ to $10/17$ and scaling to larger problems with little performance impact.
We communicate via CUDA-aware MPI and GPUDirect, avoiding the need for staging in host memory. We allocate the send and receive buffers on the GPU using OpenACC capture to guide MPI in selecting the GPU path. This strategy enables separate memory for the send/receive halo buffers.
(a) FP16/FP32
(b) FP32
(c) FP64
(d) FP64 (Baseline)
Figure 5. Visualization of a three-engine configuration and its plumes using (a) FP16, (b) FP32, (c) FP64 storage, and (d) the baseline numerics. The contours indicate where the velocity exceeds the free stream flow, and the initial state is seeded with smooth, random noise in all cases. FP32 and FP64 yield visually indistinguishable results. Visual differences in the FP16 case are solely due to the earlier onset of physical flow instabilities, yet they remain faithfully representative of the flow features. The grid-dependent nature of the baseline shock-capturing approach results in spurious grid-alignment artifacts. :
5.6. Mixed Precision
This strategy is numerically viable for various compressible flow simulations, as discussed in section˜4.3. This strategy further doubles the maximum simulation size compared to state-of-the-art shock capturing methods we are comparing against, which are numerically unstable even at single precision.
6. How Performance was Measured
6.1. HPC platforms
@@
We performed calculations on LLNL El Capitan, OLCF Frontier, and CSCS Alps. Table˜2 shows the performance attributes of these systems.
6.1.1. LLNL El Capitan
@@
The El Capitan nodes have four AMD MI300A APUs. Each APU combines 24 EPYC Zen 4 CPU cores and 228 CDNA 3 compute units with a single $128\text{\,}\mathrm{GB}\text{/}$ layer of HBM3 memory for both processor types. Nodes are interconnected via HPE Cray Slingshot-11 Ethernet in a dragonfly topology with four $200\text{\,}\mathrm{GB}\text{/}\mathrm{s}$ NICs per node. The GPUs can be programmed with unified shared memory (USM) to avoid duplicate host–device memory addresses. We use this strategy via OpenMP target offload (see section˜5.5).
6.1.2. OLCF Frontier
@@
Each Frontier node has one 64-core AMD EPYC Trento CPU and four AMD MI250X GPUs. Each MI250X GPU contains two Graphics Compute Dies (GCDs) with $64\text{\,}\mathrm{GB}\text{/}$ of HBM2E each (or $128\text{\,}\mathrm{GB}\text{/}$ per GPU). The total system memory is $9.6\text{\,}\mathrm{PB}\text{/}$, equally split between HBM2E GPU and DDR4 CPU memory. Frontier uses HPE Slingshot Ethernet for interconnects with a 3-hop Dragonfly topology and four $200\text{\,}\mathrm{GB}\text{/}\mathrm{s}$ NICs per node, which are attached to the GPUs.
6.1.3. CSCS Alps
@@
Alps nodes have 4 NVIDIA Grace Hopper GH200 superchips. Each GH200 has a Hopper GPU with $4\text{\,}\mathrm{TB}\text{/}\mathrm{s}$ of memory bandwidth for its $96\text{\,}\mathrm{GB}\text{/}$ of HBM3 memory and a Grace CPU with 72 Arm-v9 cores with $500\text{\,}\mathrm{GB}\text{/}\mathrm{s}$ bandwidth for its $120\text{\,}\mathrm{GB}\text{/}$ of LPDDR5 memory. NVLink-C2C connects the CPU and GPU, enabling efficient data movement. We use this new capability herein (see section˜5.5). Alps’s nodes interconnect through HPE Slingshot with $200\text{\,}\mathrm{GB}\text{/}\mathrm{s}$ injection bandwidth per superchip.
6.2. Software environment and performance baseline
@@
We base our implementation on MFC (bryngelson19_cpc; wilfong252) [^1] Openly available at https://github.com/MFlowCode/MFC, a compressible flow solver that implements state-of-the-art numerical shock capturing, as well as the IGR implementation of this work. All of MFC’s schemes scale ideally to 100% of LLNL El Capitan, OLCF Frontier, and CSCS Alps, among other previous and existing flagship supercomputers (radhakrishnan24; elwasif23).
@@
MFC offloads computation onto GPU and superchip/APU devices via either OpenMP or OpenACC, and uses metaprogramming to abstract away and automate vendor-specific or otherwise burdensome optimizations. MFC has a history of being used to simulate compressible multi-species, phase, and chemically reacting fluid flows (charalampopoulos21; bryngelson19_whales; bryngelson23; cisneros25). Performance results are measured using a representative three-dimensional simulation of the exhaust plume of a single Mach 10 jet.
@@
We use MFC’s optimized implementation of WENO nonlinear reconstructions and HLLC approximate Riemann solves as a baseline for performance comparisons (wilfong252). This implementation matches or outperforms other codebases with the same and similar reconstruction schemes. On Frontier and El Capitan we test HPE’s CCE 19.0.0 compiler and a pre-release of AMD’s Next Generation Flang compiler with HPE Cray MPICH-8.1.31 for messaging. For both compilers, we specify -O3 for optimization. With CCE, we also apply the -haggress flag to improve performance in large kernels. On Alps, we use a pre-release of the NVIDIA HPC SDK 25.9 and HPE Cray MPICH 8.1.30 for messaging. We also use the compile-time flag -gpu=fastmath, which yields a performance improvement for single-precision computation on NVIDIA devices. The latest NVIDIA HPC SDK exhibits an FP64 performance regression; we use NVHPC SDK 24.3 instead.
6.3. Measurement tools
@@
We measure execution time using application internal timers, including the standard cpu_time and system_clock procedures. On Frontier and El Capitan, we periodically sample the instantaneous GPU or APU power draw by reading the virtual file system for the AMD driver. The power reading is the same as reported by rocm-smi, but reading from the virtual filesystem has lower overhead. On Frontier, rocm-smi collects GPU and HBM power only. On El Capitan rocm-smi includes CPU, GPU, and memory power. On Alps, nvidia-smi records the module (CPU, GPU, and memory) power draw.
@@
On all platforms, we post-process the results to account only for power draw during time-stepping, which is then averaged and multiplied by the average time per time step. The number of grid points further normalizes energy use.
7. Performance Results
@@
Table 3. Wall time for representative simulations, quantified via nanoseconds per grid cell per time step. The baseline method is compared to the current work. For the AMD MI300A, USM mode is used, so no in-core quantities are presented; other unified cases use UVM. AMD device cases were run using Cray CCE 19.0.0 and AMD Flang Preview 7.0.5 compilers, using OpenACC and OpenMP offloading, with the best results presented. AMD Flang is the only current compiler option for native FP16/32 on all AMD GPU/APU devices. NVHPC is used for the GH200 cases.
@@
Figure 6. Weak scaling performance for a representative thruster problem on LLNL El Capitan (number of MI300As), OLCF Frontier (number of MI250Xs), and CSCS Alps (number of GH200s). The full system is indicated for each case. All runs use unified memory, FP16/32 mixed precision, and a 16 node configuration for scaling comparison, which ensures that all MPI communication directions are touched. On El Capitan, we see $\bm{97\%}$ efficiency out to 10750 MI300As. On Alps, ideal ($\bm{\approx 100\%}$) scaling is observed at 9.2K GH200s. On Frontier, we also achieve ideal ($\bm{\approx 100\%}$) efficiency at 37.6K MI250X GPUs, with a maximum problem size of 200T grid points. :
Figure 7. Strong scaling on all systems for the same problem configuration as fig.˜6. All results are for FP16/32 mixed precision, though FP32 and FP64 show similar results (FP32 shown in fig.˜8). The full system is shown for each case. Following fig.˜6, we base all speedups on an 8 node configuration, such that all communication directions are touched. :
7.1. Time step cost
Table˜3 shows the normalized grind times for a problem solved using WENO reconstructions and HLLC approximate Riemann solves with in-core computation (current state of the art, optimized implementation in MFC (wilfong252)), IGR with in-core computation, and IGR with unified memory on one GH200 on Alps, one MI250X GCD on Frontier, and one MI300A APU on El Capitan. The grind time is defined as nanoseconds per grid cell per time step, used to normalize against the different problem sizes that fit within device memory. Smaller grind times indicate shorter time to solution.
GH200 metrics are collected using separate memory mode for in-core cases and unified memory mode for out-of-core/unified cases. Performance impacts of less than 5% are observed when moving from in-core to unified computation on the GH200 architecture. We used the NVHPC 24.3 compiler for the GH200 FP64 cases due to performance regressions in FP64 performance in newer versions. However, an unrelated regression in NVHPC 24.3 was observed when switching from separate memory to unified memory in-core computation, resulting in a $9\%$ performance hit in that case (see table˜3). The performance difference between in- and out-of-core unified memory builds is less than $5\%$.
@@ 7.1. Time step cost
The increased grind time in calculations that use unified memory to increase problem size per device on Frontier and Alps results from the exchange of conservative variable buffers between the CPU and GPU at each Runge–Kutta update. On all devices, the time to solution is reduced by a factor of approximately 4 when comparing WENO to IGR in FP64. FP64 is broadly recommended for ENO-type shock-capturing due to catastrophic cancellation (Brogi2024OpenFOAMPrecision). Our approach can even handle even mixed FP16/FP32 precision. This reduces the time to solution by a factor of at least 6 compared to the baseline.
@@ 7.2. Scaling
We exceed 200T grid cells, or 1 quadrillion degrees of freedom (5 state variables) on Frontier using 37.6K MI250X GPUs (9408 nodes), accommodating $1386^{3}$ grid points per GCD with UVM, and FP16/32 mixed precision. On Alps, a problem size of $1611^{3}$ is used per GH200 and UVM with FP16/32 mixed precision, amounting to 45T grid point simulation on the full system (2688 nodes). On the recently deployed JSC JUPITER, this amounts to 100.3T grid points or 501T degrees of freedom, given its size and matching architecture to Alps. On El Capitan, we use $1380^{3}$ grid points per MI300A GPU and reach 113T grid points using 10750 nodes (11.1K nodes full system). The lower grid count on Alps and El Capitan is due purely to system size.
Refer to caption :
Figure 8. Strong scaling following fig.˜7, here for FP32 on Frontier. We compare the performance of the current work to the optimized baseline method as labeled. The current work uses and can accommodate 10.5B grid points per node, while the baseline accommodates 421M grid points, both of which are used for the 8-node speedup reference.
Strong scaling results on Alps, Frontier, and El Capitan are shown in fig.˜7. The base case uses $8$ nodes ($32$ GPUs), resulting in $32$ ranks on El Capitan and Alps and $64$ ranks on Frontier, arranged in a rectilinear configuration. We use FP16/32 mixed precision for storage/computation; unified memory is enabled via UVM on Frontier and Alps and USM on El Capitan; GPU-aware MPI is used on Alps. For a 32-fold increase in device count, we achieve strong scaling efficiencies of 90%, 90%, and 86% on El Capitan, Frontier, and Alps. When scaling to the full systems, we achieve 44% (El Capitan), 44% (Frontier), and 80% (Alps) strong scaling efficiencies. Thus, one can execute an 8 node computation on the full system, decreasing time to solution by a factor of about 500.
@@
7.3. Energy efficiency
Table˜4 shows the measured energy consumption per grid cell per time step, comparing our IGR implementation with a prior state-of-the-art WENO implementation. The problem size is adjusted to exhaust the total GPU/APU memory on a single node in double precision. On the NIVIDA GH200 we measured the in-core implementation. On the AMD MI250X and MI300A, we use the CCE 19.0.0 compiler with GPU-only (in-core) and unified memory builds.
The IGR method significantly reduces energy consumed compared to the current state of the art on all platforms. The most significant reduction comes from the improved time to solution, and to lowest order, the Frontier and El Capitan energy improvements match the grind time speedups in table˜3. Due to the higher power draw of the WENO scheme on Alps, we observe energy savings beyond those resulting from grind time speedup.
The largest improvement is realized on Frontier with a $5.38$-times improvement in energy consumed by the GPU. As described in section˜6.3, the Frontier measurements include only GPU and GPU HBM energy and do not represent changes that may have impacted energy consumption of the CPU or CPU memory.